Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
boldrini, capocasa - 23:32 Monday 04 July 2022 (56346) Print this report
ISC shift: ITF recovery after TCS actuators alignment

The goal of the shift was to recover the lock acquisition after the re-alignment of the TCS actuators.

For most of our shift we waited for TCS staff to complete the alignment of their actuators. Once they were satisfied, we recovered the lock of the arms and performed a scan of the cavities to measure the mismatch in this configuration (Fig.1).

They are both around 2%. North arm is better with respect to the last measurement.

We then proceeded with the lock of the DRMI. This was an extremely difficult task, as we could only obtain some short locks after changing the demodulation phase of MICH/SRCL by pi (2.1 -> -1.0 rad) and after increasing the SRCL gain from 0.4 to 0.6. The sideband power was also very low, and to increase the density of the triggers we modified the tresholds:

  • b4_dc_max = 20
  • b4_12_max = 0.03
  • b4_112_max = 0.015

Even with these changes, motivated by observing the error signals and corrections on the triggers, it was impossible to lock the CITF without misaligning the SR by at least 4 urad. After the ITF locked, we started slowly re-aligning the SR towards the position found with the pre-alignmnent, but this broke the lock (Fig.2). Even so, the sidebands are noticeably lower than usual (about half the usual power at this point). We don't suffer from a saturation issue on SRCL as in the last CITF lock attempt.
The image from the phase-cameras taken during this lock does not paint a better picture (Fig.3), although the SR was misaligned (2 urad) during this lock.

We could not re-lock the ITF to try and adjust the MICH/SRCL demodulation phase while keeping the SR misaligned.

Before concluding the shift, we re-measured the mismatch in the arms (Fig.4), and found a better matching for west arm after the thermalization finished (1.7%)

We leave the arms locked on the IR.

Images attached to this report
Comments to this report:
boldrini - 15:54 Tuesday 05 July 2022 (56356) Print this report

We reverted the changes in the DRMI.ini file. The parameters are the same used before the re-alignment of the TCS actuators.

Search Help